OpenStudy (anonymous):

in my book it says to simplify this 6(8+3c)+4(10+2c) how do i solve this ?

OpenStudy (anonymous):

expand and simplify\[6(8+3c)+4(10+2c)=48+18c+40+8c=26c+88\]

OpenStudy (anonymous):

i know but how?

OpenStudy (anonymous):

srry never mind got it :P

OpenStudy (anonymous):

:-D

OpenStudy (anonymous):

is there a difference with multiplication ?